OpenStudy (anonymous):

Solve: | m + 8| = 16 {8} {-24, 8} {8, 16} {-24}

OpenStudy (anonymous):

m=8,-24

OpenStudy (cwrw238):

this generates 2 equations m + 8 = 16 and m + 8 = -16 solve these to get the solution set

OpenStudy (cwrw238):

K-K-K-Kiaa - plz dont just give the answer - we are here to show people how to do the problem